MatLab Problem: Please use matlab format and follow instruction on the problem, thank you in advance!

Problem 2 (1 point. Instructors Initials) Determine all values of x that satisfy the following system of non-linear equations: y = sin(x) Graph the two functions, and hand-write the solution (or solutions, if more than one) on the graph. Include your name in the title of the graph.
Answer #1

MATLAB Codes for Plot Nonlinear Functions % Plot of four nonlinear func iotns % Function funcA subplot (2,2,1) X-2:0.2:2 Y-X.y=x43x-1 y-e-sin(x) 0.5 y-e-sinp) О х -3.1831 -0.5 -10 O x=0.322 -3.5 -2.5 y-e-1.5-tanx y=340.5*sin (x)-x 4 2 一yses_1.5-tan-1